A parallel plate capacitor of 2 pF is charged to 12 V and then isolated from the power source. The space between the plates is now filled with a dielectric plate sheet whose dielectric constant is 5.0. The new voltage across the capacitor is

with steps please
Answer #1

Given C = 2pF = 2*10-12 F

V =12 V

First Find the charge

Q =CV

Q = 2*10-12*12 = 2.4 *10-11C

Q= 2.4*10-11 A

-------------

When a dielectric is inserted the capaitance increases

Cnew = kC = 5*C = 10pF

Charge Q remains same ,so new value of voltage across the capacitor is

V =Q/C = 2.4*10-11/10*10-12

V =2.4 V

ANSWER: V =2.4 V

---------------------